5000 Joules of heat are added to 100 grams of water at 30 Celsius. Calculate the final temperature. Cpwater = 4.184 J/g C
Answer #1

Answer:

We know that heat q=mCp(Tf-Ti)

Where q= heat, m=mass, Cp=specific heat, Ti=initial temperature and Tf=final temperature.

Given q=5000J, m=100g, Cp=4.184 J/g°C, Ti=30°C, Tf=?

Tf-Ti=q/mCp

Tf=Ti+(q/mCp)

Tf=30°C + (5000 J/100 g x 4.184 J/g°C)

Tf=30°C + 11.95 °C

Tf=41.95 °C

Thanks and I hope you like it.